OpenStudy (anonymous):

The ratio of the height of the pyramid to the edge of the square base is 1.5 to 1. The height of the pyramid is 3 meters. What is the approximate length of the slant height of the pyramid?

OpenStudy (anonymous):

A) 4.4 m B) 3.2 m C) 2.8 m D) 1.4 m SHOW ALL WORK.

OpenStudy (anonymous):

when you say the edge of the square base you mean the length of one side of the square base - right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

length of square edge = 3 / 1.5 = 2 distance of corner of square from middle of square is by pythagors sqrt(1^2 + 1^2) = sqrt 2 consider the right angled triangle formed by perpendicular height, distance from center and slant height, by pythagoras h^2 = 3^2 +( sqrt2)^2 = 9 + 2 = 11 = sqrt 11 = approx 3.2 b)

OpenStudy (anonymous):

ok with that?